OpenStudy (anonymous):

HELP PLEASE? Area = x^2 + 40x + 300x Find the dimensions ( l x w) and the area of the foundation if x = 20 ft. I already solved x + 30 and x + 10 but can't figure out where to put the 20? When I replace it with 40x I get no results!

OpenStudy (anonymous):

Is it x^2+40x+300???

OpenStudy (anonymous):

@loopgirl

OpenStudy (anonymous):

That is my original question. Then on the second one it says x + 20.

OpenStudy (anonymous):

Oh you wrote 300x in this question However if it is 300, then x+10 and x+30 are dimensions , putting x=20 you get the dimensions as 30 and 50 and hence the area is 30*50=1500

OpenStudy (anonymous):

feet is the unit. after computing the result, you have to write area=1500 sq. feet

OpenStudy (anonymous):

Oh my bad! So the equation would be wrote as x^2 + 20x + 300 x + 30 = x + 50?

OpenStudy (anonymous):

no question is absolutely correct. it is x^2+40x+300

OpenStudy (anonymous):

I'm confused. :( How does x = 20 work into the equation?

OpenStudy (anonymous):

You just put x=20 for the solution of the last part. You are not asked to put x=20 in the first question. 1st question's solution is length=x+30 and breadth =x+10

OpenStudy (anonymous):

And you put x=20, then (20)^2+40*20+300=400+800+300=1500sq feet the answer of the last question

OpenStudy (anonymous):

OH. So it would be 40 * 20 and then I solved for that, which you just did. Thank you so much, I get it now!!

OpenStudy (anonymous):

Yeah...